OpenStudy (anonymous):

Please help me solve this equation: write in exponential form sqrt (7) x^3y^5z^2 WITH SOLUTION tnx!!!

OpenStudy (allank):

Most of the equation is already in exponent form. Simply note that \[\sqrt{a}= a^(0.5)\]

OpenStudy (anonymous):

\[\sqrt{a }=a^{1/2}\] use thie rule

OpenStudy (anonymous):

\[\sqrt{7x^3y ^{5}z ^{2}}=(7x^3y^5z^2)^{1/2}\] @ynapot97 : hope this will help you